### v3

# Re-Reading Improves Reasoning in Large Language Models

Abstract
To enhance the reasoning capabilities of off-the-shelf Large Language Models (LLMs), we introduce a simple, yet general and effective prompting method, Re2, i.e., Re-Reading the question as input. Unlike most thought-eliciting prompting methods, such as Chain-of-Thought (CoT), which aim to elicit the reasoning process in the output, Re2 shifts the focus to the input by processing questions twice, thereby enhancing the understanding process. Consequently, Re2 demonstrates strong generality and compatibility with most thought-eliciting prompting methods, including CoT. Crucially, Re2 facilitates a "bidirectional" encoding in unidirectional decoder-only LLMs because the first pass could provide global information for the second pass. We begin with a preliminary empirical study as the foundation of Re2, illustrating its potential to enable "bidirectional" attention mechanisms. We then evaluate Re2 on extensive reasoning benchmarks across 14 datasets, spanning 112 experiments, to validate its effectiveness and generality. Our findings indicate that, with the exception of a few scenarios on vanilla ChatGPT, Re2 consistently enhances the reasoning performance of LLMs through a simple re-reading strategy. Further analyses reveal Re2’s adaptability, showing how it can be effectively integrated with different LLMs, thought-eliciting prompting, and ensemble strategies.111  Our code is available at https://github.com/Tebmer/Rereading-LLM-Reasoning/
Re-Reading Improves Reasoning in Large Language Models
Xiaohan Xu1thanks:   This work was done during internship at Microsoft., Chongyang Tao2, Tao Shen3, Can Xu2, Hongbo Xu1, Guodong Long3, Jian-guang Lou2 1Institute of Information Engineering, CAS, {xuxiaohan,hbxu}@iie.ac.cn 2Microsoft Corporation, {chotao,caxu,jlou}@microsoft.com 3AAII, School of CS, FEIT, UTS, {tao.shen,guodong.long}@uts.edu.au
1 Introduction
In the ever-evolving landscape of artificial intelligence, large language models (LLMs) have emerged as a cornerstone of natural language understanding and generation (Brown et al., 2020; Touvron et al., 2023a, a; OpenAI, 2023). As these LLMs have grown in capability, a pivotal challenge has come to the forefront: imbuing them with the ability to reason effectively. The capacity to engage in sound reasoning is a hallmark of human intelligence, enabling us to infer, deduce, and solve problems. In LLMs, this skill is paramount for enhancing their practical utility. Despite their remarkable capabilities, LLMs often struggle with nuanced reasoning (Blair-Stanek et al., 2023; Arkoudas, 2023), prompting researchers to explore innovative strategies to bolster their reasoning prowess (Wei et al., 2022b; Gao et al., 2023; Besta et al., 2023).
Figure 1: Example inputs of CoT prompting versus CoT prompting with Re2. Re2 is a simple prompting method that repeats the question as input. Typically, tokens in the question, such as "tennis balls", cannot see subsequent tokens in the original setup for LLMs (the top figure). In contrast, LLMs with Re2 allows "tennis balls" in the second pass to see the entire question containing "How many …", achieving an effect of a "bidirectional" understanding (the bottom figure).
Figure 2: Illustration of the attention distribution of each token in the second pass attending to the entire input in LLaMA-2 using the GSM8K dataset. A darker cell indicates higher attention. The region within the dashed upper triangle demonstrates that every token in the second pass has obvious attention to its subsequent tokens in the first pass. This suggests that re-reading in LLMs is promising for achieving a "bidirectional" understanding of the question.
Existing research on reasoning has predominantly concentrated on designing diverse thought-eliciting prompting strategies to elicit reasoning processes in the output phase, such as Chain-of-Thought (CoT) Wei et al. (2022b), Program-Aided Language Model (PAL) Gao et al. (2023), etc. Yao et al. (2023a); Besta et al. (2023); Wang et al. (2023a). In contrast, scant attention has been paid to the understanding of the input phase. In fact, comprehension is the first step before solving the problem, which is crucially important. However, in the era of generative AI, most LLMs adopt the decoder-only LLMs with unidirectional attention, like GPT-3 Brown et al. (2020) and LLaMA Touvron et al. (2023b). Compared with encoder-based language models that feature bidirectional attention Devlin et al. (2019), the unidirectional attention limits every token’s visibility to only previous tokens when encoding a question, potentially impairing the global understanding of the question Du et al. (2022) (the top figure in Figure 1). Fortunately, many cognitive science studies have revealed that humans tend to re-read questions during learning and problem-solving to enhance the comprehension process Dowhower (1987, 1989); Ozek and Civelek (2006). Motivated by this, we also conduct a preliminary empirical study for LLaMA-2 Touvron et al. (2023b) by repeating the question two times using the GSM8K dataset Cobbe et al. (2021). Figure 2 shows that LLaMA-2 with re-reading is promising to achieve a "bidirectional" understanding of the question, and further improve the reasoning performance.
Based on the observation and inspired by the human strategy of re-reading, we present a simple yet effective and general reasoning prompting strategy, Re2, i.e., Re-Reading the question as input (see the illustration in Figure 1). Similar to the human problem-solving process, where the primary task is to comprehend the problem, we place our emphasis on designing the prompting strategy on the input phase. Thus, Re2 is orthogonal to and compatible with most thought-eliciting prompting methods in the output phase, such as CoT, PAL, etc. Moreover, instead of processing the input in only a single pass, the repetition of questions enables LLMs to allocate more computational resources to input encoding, akin to "horizontally" increasing the depth of neural networks. Therefore, LLMs with Re2 is promising to have a deeper understanding of the question and improve reasoning performance. More interestingly, LLMs with Re2 show potential for a "bidirectional" understanding of questions in the context of unidirectional LLMs. This is because every token in the second pass could also attend to its subsequent tokens in the first pass (see illustration in Figure 1 and Figure 2).
To validate the efficacy and generality of Re2, we conducted extensive experiments spanning arithmetic, commonsense, and symbolic reasoning tasks across 14 datasets and 112 experiments. The results show that, with the exception of certain scenarios on vanilla ChatGPT, our Re2 with a simple re-reading strategy consistently enhances the reasoning performance of LLMs. Re2 exhibits versatility across various LLMs, such as Text-Davinci-003, ChatGPT, LLaMA-2-13B, and LLaMA-2-70B, spanning both instruction fine-tuning (IFT) and non-IFT models. We also explore Re2 in task settings of zero-shot and few-shot, thought-eliciting prompting methods, and the self-consistency setting, highlighting its generality.
2 Methodology
2.1 Vanilla Chain-of-Thought for Reasoning
We begin with a unified formulation to leverage LLMs with CoT prompting to solve reasoning tasks. In formal, given an input $x$ and a target $y$, a LLM $p$ with CoT prompting can be formulated as
$y\sim\sum\limits_{z\sim~{}p({\textnormal{z}}|C_{x})}p({\textnormal{y}}|C_{x},z)\cdot p(z|C_{x}),$
where $\text{where }C_{x}=\operatorname*{c}\nolimits^{\text{(cot)}}(x).$ (1)
In this formulation, $C_{x}$ denotes the prompted input. $\operatorname*{c}^{(\text{cot})}(\cdot)$ represents the template with CoT prompting instructions, such as ‘let’s think step by step’. z stands for a latent variable of rationale, and $z$ denotes a sampled rationale in natural language. Consequently, the LLMs can break down complex tasks into more manageable reasoning steps, treating each step as a component of the overall solution chain. We employ CoT as a baseline to solve reasoning tasks without compromising its generality. In addition to CoT, our proposed simple Re2 can serve as a "plug & play" module adaptable to most other prompting methods (§2.3).
2.2 Re-Reading (Re2) Improves Reasoning
Drawing inspiration from the human strategy of re-reading, we introduce this strategy for LLM reasoning, dubbed Re2, to enhance understanding in the input phase. With Re2, the prompting process in Eq. 1 can be readily rephrased as:
$y\sim\sum_{z\sim~{}p({\textnormal{z}}|C_{x})}p({\textnormal{y}}|C_{x},z)\cdot p(z|C_{x}),$
where $\text{where }C_{x}=\operatorname*{c}\nolimits^{(\text{cot})}(\operatorname*{re2}(x)).$ (2)
[…]
Re2 Prompting Q: {Input Query} Read the question again: {Input Query} # Thought-eliciting prompt (e.g.,‘‘Let’s think step by step") #
where ‘{Input Query}‘ is a placeholder for the input query, $x$. The left part of this prompting could incorporate other thought-eliciting prompts. Intuitively, Re2 offers two advantages to enhance the understanding process: (1) it allocates more computational resources to the input, and (2) it facilitates a "bidirectional" understanding of the question, where the first pass provides global information for the second pass.
2.3 Generality of Re2
Due to Re2’s simplicity and emphasis on the input phase, it can be seamlessly integrated with a wide range of LLMs and algorithms, including few-shot settings, self-consistency, various thought-eliciting prompting strategies, and more. We offer insights into the integration of Re2 with other thought-eliciting prompting strategies as an illustration.
Compared with those thought-eliciting prompting strategies that focus on the output phase, Re2 shifts the emphasis towards understanding the input. Therefore, Re2 exhibits significant compatibility with them, acting as a “plug & play” module. This synergy has the potential to further enhance the reasoning abilities of LLMs. With a specific thought-eliciting prompting, $\tau$, designed to elicit thoughts from the LLMs, Eq. (2.3) is rewritten as:
$y\sim\sum_{z\sim~{}p({\textnormal{z}}|C_{x})}p({\textnormal{y}}|C_{x},z)\cdot p(z|C_{x}),$
where $\text{where }C_{x}=\operatorname*{c}\nolimits^{(\tau)}(\operatorname*{re2}(x)).$ (3)
Here, $\tau$ denotes various thought-eliciting promptings beyond CoT, such as Plan-and-Solve Wang et al. (2023a), and Program-Aided Prompt Gao et al. (2023), etc. We also conducted lots of experiments to validate the generality of Re2 in §3.4.
3 Experiments
LLMs Methods GSM SVAMP ASDIV AQUA MultiArith SingleEQ AddSub
davinci-003 Vanilla 19.48 67.60 69.00 28.74 31.33 86.22 89.87
Vanilla+Re2 24.79 $\uparrow$ 5.31 70.90 $\uparrow$ 3.30 71.20 $\uparrow$ 2.20 30.31 $\uparrow$ 1.57 42.33 $\uparrow$ 11.00 87.20 $\uparrow$ 0.98 92.15 $\uparrow$ 2.28
CoT 58.98 78.30 77.60 40.55 89.33 92.32 91.39
CoT+Re2 61.64 $\uparrow$ 2.68 81.00 $\uparrow$ 2.70 78.60 $\uparrow$ 1.00 44.49 $\uparrow$ 3.94 93.33 $\uparrow$ 4.00 93.31 $\uparrow$ 0.99 91.65 $\uparrow$ 0.26
ChatGPT Vanilla 77.79 81.50 87.00 63.39 97.83 95.28 92.41
Vanilla+Re2 79.45 $\uparrow$ 1.66 84.20 $\uparrow$ 2.70 88.40 $\uparrow$ 1.40 58.27 $\downarrow$ 5.12 96.67 $\downarrow$ 1.16 94.49 $\downarrow$ 0.79 91.65 $\downarrow$ 0.76
CoT 78.77 78.70 85.60 55.91 95.50 93.70 88.61
CoT+Re2 80.59 $\uparrow$ 1.82 80.00 $\uparrow$ 1.30 86.00 $\uparrow$ 0.40 59.06 $\uparrow$ 3.15 96.50 $\uparrow$ 1.00 95.28 $\uparrow$ 1.58 89.87 $\uparrow$ 1.26
Table 1: Results on arithmetic reasoning benchmarks. denotes that Vanilla is even superior to CoT prompting.
LLMs Methods Commonsense Symbolic
CSQA StrategyQA ARC-e ARC-c ARC-t Date Coin
davinci-003 Vanilla 74.20 59.74 84.81 72.01 80.58 40.92 49.80
Vanilla+Re2 76.99 $\uparrow$ 2.79 59.91 $\uparrow$ 0.17 88.22 $\uparrow$ 3.41 75.68 $\uparrow$ 3.67 84.07 $\uparrow$ 3.49 42.01 $\uparrow$ 1.09 52.40 $\uparrow$ 2.60
CoT 71.66 67.55 85.69 73.21 81.57 46.07 95.60
CoT+Re2 73.05 $\uparrow$ 1.39 66.24 $\downarrow$ 1.31 87.84 $\uparrow$ 2.15 76.02 $\uparrow$ 2.81 83.94 $\uparrow$ 2.37 52.57 $\uparrow$ 6.50 99.60 $\uparrow$ 4.00
ChatGPT Vanilla 76.66 62.36 94.32 85.41 91.37 47.43 52.00
Vanilla+Re2 78.38 $\uparrow$ 1.72 66.99 $\uparrow$ 4.63 93.81 $\downarrow$ 0.51 83.19 $\downarrow$ 2.22 90.30 $\downarrow$ 1.07 47.97 $\uparrow$ 0.54 57.20 $\uparrow$ 5.20
CoT 69.94 67.82 93.35 83.53 90.11 43.63 88.80
CoT+Re2 71.66 $\uparrow$ 1.72 69.34 $\uparrow$ 1.52 93.14 $\downarrow$ 0.21 84.47 $\uparrow$ 0.94 90.27 $\uparrow$ 0.16 47.15 $\uparrow$ 3.52 95.20 $\uparrow$ 6.40
Table 2: Results on commonsense and symbolic reasoning benchmarks. denotes that Vanilla is even superior to CoT prompting.
3.1 Benchmarks
We assess Re2 prompting across three key categories of reasoning benchmarks. Details of all datasets are shown in Appendix A
Arithmetic Reasoning
We consider the following seven arithmetic reasoning benchmarks: the GSM8K benchmark of math word problems Cobbe et al. (2021), the SVAMP dataset of math word problems with varying structures Patel et al. (2021), the ASDiv dataset of diverse math word problems Miao et al. (2020), the AQuA dataset of algebraic word problems Ling et al. (2017), the AddSub  Hosseini et al. (2014) of math word problems on addition and subtraction for third to fifth grader, MultiArith Roy and Roth (2015) dataset of math problems with multiple steps, and the SingelEQ Roy et al. (2015) dataset of elementary math word problems with single operation.
Commonsense and Symbolic Reasoning
For commonsense reasoning, we use CSQA Talmor et al. (2019), StrategyQA Geva et al. (2021), and the ARC Clark et al. (2018). CSQA dataset consists of questions that necessitate various commonsense knowledge. The StrategyQA dataset comprises questions that demand multi-step reasoning. The ARC dataset (denoted as ARC-t) is divided into two sets: a Challenge Set (denoted as ARC-c), containing questions that both retrieval-based and word co-occurrence algorithms answered incorrectly, and an Easy Set (denoted as ARC-e). We evaluate two symbolic reasoning tasks: date understanding Suzgun et al. (2023a) and Coinflip Wei et al. (2022b). Date understanding is a subset of BigBench datasets Suzgun et al. (2023a), which have posed challenges for previous fine-tuning efforts. Coinflip is a dataset of questions on whether a coin is still heads up after it is flipped or not based on steps given in the questions.
3.2 Language Models and Implementations
Baseline Prompting. In our implementation, we rigorously evaluate the performance of our Re2 model on two baseline prompting methods: Vanilla and CoT. The Vanilla approach aligns with the standard prompting method outlined in Wei et al. (2022b); Kojima et al. (2022), wherein no specific prompts are employed to elicit thoughts from LLMs. Conversely, the CoT method guides the model through a step-by-step thought process.
Re2 Prompting. We incorporate our Re2 strategy into these baseline methods to assess its impact, denoted as Vanilla+Re2 and CoT+Re2. To avoid the impact of randomness introduced by the demonstrations in a few-shot setting, we mainly assess our method in a zero-shot setting, following Chen et al. (2023); Wang et al. (2023a); Du et al. (2023). Additionally, for different tasks, we design answer-format instructions in prompts to regulate the structure of the final answer, facilitating precise answer extraction. Detailed information regarding the baseline prompting, Re2 prompting, and answer-format instructions can be found in the paper’s Appendix B.
Implementations. Our decoding strategy uses greedy decoding with a temperature setting of 0, thus leading to deterministic outputs. For these experiments, we employ two powerful backbones: ChatGPT (gpt-3.5-turbo-0613) OpenAI (2022) and davinci-003 (text-davinci-003) [^2] https://platform.openai.com/docs/models/gpt-3-5 , across all prompting methods, including Vanilla, CoT, Vanilla+Re2, and CoT+Re2.
3.3 Evaluation Results
Table 1 presents the results on arithmetic reasoning datasets, and Table 2 on commonsense reasoning and symbolic reasoning. In almost all scenarios, LLMs with Re2 achieve consistent improvements across both LLMs (davinci-003 and ChatGPT) and prompting methods (Vanilla and CoT). Specifically, davinci-003 with Vanilla+Re2 shows average improvements of 3.81, 2.51, and 1.85 in arithmetic, commonsense, and symbolic tasks, respectively. With CoT, davinci-003 generates intermediate reasoning steps, significantly enhancing the reasoning performance of LLMs. By applying Re2, davinci-003 with CoT+Re2 demonstrates further improvement, with average gains of 2.22, 1.23, and 5.25 in the same categories, respectively. These results indicate that Re2 can benefit LLMs in directly generating answers and improve the performance of CoT leading to correct answers.
When applied to ChatGPT, Re2 exhibits consistent improvement on most datasets, except for a slight drop in performance on a few datasets, e.g., AQUA and MultiArith, when using Vanilla+Re2. This exception could be due to ChatGPT’s exposure to these datasets with CoT outputs during instruction fine-tuning (IFT) Chen et al. (2023). On such datasets, ChatGPT with Vanilla still produces CoT-like output (see examples in Appendix E) and even outperforms ChatGPT with CoT (as indicated by the results in Tables 1 and 2). Chen et al. (2023) obtained similar experimental results and suggested that this occurs because ChatGPT may have been exposed to these task datasets containing CoT explanations without explicit prompting. Therefore, additional explicit instructions, like CoT or Re2, might disrupt this learned pattern in ChatGPT, possibly leading to decreased performance. Nonetheless, on some datasets like SVAMP, ASDIV, CSQA, and Date, Re2 still manages to improve the baseline Vanilla prompting. Moreover, in datasets where CoT prompting normally surpasses Vanilla prompting, such as GSM, StrategyQA, and Coin, Re2 significantly enhances Vanilla prompting ($\uparrow$ 4.63 on StrategyQA and $\uparrow$ 5.20 on the Coin dataset). Overall, our Re2 method still achieves improvements in 71% of the experiments on ChatGPT. More examples from the experiment results can be found in Appendix E.
Figure 3: Evaluation results of the times of reading on GSM benchmark.
3.4 Discussions
Times of Question Reading
We delve deeper into the impact of the times of question re-reading on reasoning performance. Figure 3 illustrates how the performance of two distinct LLMs evolves concerning various times of question re-reading. An overarching pattern emerges across all models: performance improves until the number of re-reads reaches 2 or 3, after which it begins to decline with further increases in question re-reading times. The potential reasons for inferior performance when reading the question multiple times are two-fold: i) overly repeating questions may act as demonstrations to encourage LLMs to repeat the question rather than generate the answer, and ii) repeating the question significantly increase the inconsistency of the LLMs between our inference and pretraining/alignment (intuitively in the learning corpora, we usually repeat a question twice to emphasize the key part, rather not more). It’s noteworthy that reading the question two times tends to be optimal for accommodating most scenarios in our experiments, which is why we refer to this practice as “re-reading” in our paper.
LLMs Methods GSM
ChatGPT PS 75.59
PS+Re2 76.27
PAL 75.59
PAL + Re2 79.38
davinci-003 PS 55.65
PS+Re2 58.68
PAL 68.61
PAL + Re2 70.20
Table 3: Evaluation results of some thought-eliciting promptings beyond CoT with Re2.
Compatibility with Thought-Eliciting Prompt Strategies
Compared to previous methods attempting to elicit thoughts in the output from LLMs, our Re2 emphasizes the understanding of the input. Therefore, we are intrigued to explore whether Re2 is effective with various thought-eliciting prompting strategies other than CoT. To investigate this, we apply Re2 to two other recently introduced prompting methods, namely, Plan-and-Solve (PS) Wang et al. (2023a) and Program-Aided Language models (PAL) Gao et al. (2023). The former model devises a plan to divide the entire task into smaller subtasks, and then carries out the subtasks according to the plan, while the latter generates programs as the intermediate reasoning steps. We directly apply our Re2 to these two methods by making a simple alteration to the input by repeating the question. Table  3 presents the evaluation findings on the GSM benchmark. Our observations reveal a consistent trend, akin to what was observed with chain-of-thought prompting. These results suggest that the effectiveness of our Re2 mechanism generally extends across various prompting methodologies.
LLMs Methods GSM SVAMP ASDIV AQUA MultiArith SingleEQ AddSub
davinci-003 Vanilla 16.98 69.10 70.56 28.34 38.67 83.46 88.86
Vanilla+Re2 19.02 $\uparrow$ 2.04 73.60 $\uparrow$ 4.50 73.23 $\uparrow$ 2.67 27.95 $\downarrow$ 0.39 46.00 $\uparrow$ 7.33 84.06 $\uparrow$ 0.60 89.37 $\uparrow$ 0.51
CoT 56.63 78.90 79.96 46.45 96.16 90.94 88.60
CoT+Re2 60.12 $\uparrow$ 3.49 79.80 $\uparrow$ 0.90 81.21 $\uparrow$ 1.25 44.89 $\downarrow$ 1.56 96.83 $\uparrow$ 0.67 91.14 $\uparrow$ 0.20 89.37 $\uparrow$ 0.77
Table 4: Evaluation results on arithmetic reasoning benchmarks under few-shot setting.
LLMs Methods GSM SVAMP ASDIV AQUA MultiArith SingleEQ AddSub
Llama-2-13B Vanilla 5.76 43.90 52.91 22.44 6.33 68.11 66.58
Vanilla+Re2 6.82 $\uparrow$ 1.06 47.90 $\uparrow$ 4.00 53.15 $\uparrow$ 0.24 17.32 $\downarrow$ 5.12 6.50 $\uparrow$ 0.17 69.68 $\uparrow$ 1.57 70.12 $\uparrow$ 3.54
CoT 21.99 41.60 45.18 22.83 56.83 58.46 58.99
CoT+Re2 22.37 $\uparrow$ 0.38 46.50 $\uparrow$ 4.90 48.81 $\uparrow$ 3.63 24.80 $\uparrow$ 1.97 55.83 $\downarrow$ 0.99 66.34 $\uparrow$ 7.88 60.76 $\uparrow$ 1.77
Llama-2-70B Vanilla 11.60 56.60 61.31 20.08 24.67 77.17 80.25
Vanilla+Re2 13.50 $\uparrow$ 1.90 63.60 $\uparrow$ 7.00 64.66 $\uparrow$ 3.35 22.05 $\uparrow$ 1.97 25.00 $\uparrow$ 0.33 80.31 $\uparrow$ 3.14 84.05 $\uparrow$ 3.80
CoT 49.73 66.90 68.08 37.80 79.83 80.51 74.18
CoT+Re2 56.71 $\uparrow$ 6.98 70.40 $\uparrow$ 3.50 70.42 $\uparrow$ 2.34 38.58 $\uparrow$ 0.78 88.83 $\uparrow$ 9.00 81.10 $\uparrow$ 0.59 69.37 $\downarrow$ 4.81
Table 5: Evaluation results of LLAMA-2 on arithmetic reasoning benchmarks
Compatibility with Few-Shot Prompting
It is noteworthy that our proposed re-reading mechanism is compatible with few-shot prompting. To demonstrate this compatibility, we conducted experiments on arithmetic reasoning tasks using the davinci-003 model, employing both Vanilla and CoT prompting methods. The few-shot prompting strategy and exemplars used align with those presented in Wei et al. (2022b). For both the Vanilla+Re2 and CoT+Re2 methods, we applied the re-reading mechanism to the exemplars as well. The results of these experiments are presented in Table 4. We can observe that the inclusion of the re-reading mechanism consistently enhances the performance of both prompting methods, mirroring our findings in the zero-shot setting.
Effect on Non-IFT Models
In our primary experiments, we employed the ChatGPT and davinci-003 models, which had undergone IFT training. These models, being aligned with human-like behavior, are better equipped to follow instructions effectively. Additionally, they may have been exposed to datasets with CoT prompting during their training, making the “re-reading” mechanism potentially more beneficial in recalling explanations. To gauge the broader applicability of our approach and to eliminate any IFT-related impacts, we conducted experiments on non-IFT pretrained models: Llama-2-13B and Llama-2-70B Touvron et al. (2023b). Llama-2 is an open-source model pretrained on publicly available data without IFT or RLHF fine-tuning. We evaluated Llama-2 on arithmetic reasoning tasks under a zero-shot setting, following Kojima et al. (2022). The results are presented in Table 5. The results clearly indicate that the re-reading mechanism consistently enhances the performance of both Vanilla and CoT prompting methods across most tasks when applied to Llama-2 models. This observation underscores the generality of our approach and dispels concerns about potential data leakage from IFT during training. This also underscores the versatility of Re2, which can be effectively employed across various model scales and types, regardless of whether they have undergone IFT training or are non-IFT LLM.
LLMs Methods GSM SVAMP
ChatGPT Vanilla 77.79 81.50
Vanilla+SC 85.52 87.50
Vanilla+Re2 +SC 85.22 87.70
CoT 78.77 78.70
CoT+SC 85.75 84.90
CoT+Re2 +SC 86.88 87.70
Table 6: Evaluation results of re-reading with self-consistency on GSM benchmark.
Compatibility with Self-consistency
Existing research indicates that the chain-of-thought prompting approach can be enhanced by adopting the self-consistency method, which involves aggregating the majority final answer from multiple sampled generations. We are also intrigued by the potential for further enhancing the proposed re-reading mechanism using this method. Consequently, we conduct experiments testing the integration of Re2 with the self-consistency approach on the GSM benchmark with temperature of 0.7 when decoding (Table 6). Our findings demonstrate that self-consistency significantly enhances the performance of both prompting methods. Despite self-consistency’s aggregation of multiple answers, our re-reading mechanism still contributes to improvement on most scenarios, indicating its compatibility with the self-consistency approach.
Figure 4: Left figure: model performance versus complexity of questions. X-axis means the complexity of questions and Y-axis refers to frequency. The gray hist means the number of total cases for each complexity. Right figure: n-gram recall between the generation and the input question. We take the question and generation as the reference and hypothesis respectively.
Performance across Different Question Complexity.
We further investigate the impact of input question complexity on the reasoning performance of both CoT and CoT+Re2 promptings using ChatGPT on GSM8K dataset, as shown in the left part of Figure 4. In accordance with Fu et al. (2022), we measure question complexity by counting the reasoning steps present in the ground-truth explanations. Our findings reveal a noticeable trend: the performance of all promptings generally diminishes as question complexity increases, suggesting that the current LLMs still struggle with handling intricate queries. Notably, the introduction of re-reading enhances performance on various complexities, including those slightly complex questions. This observation underscores the benefits of Re2 for improving reasoning capabilities over complex questions. To further validate the improved understanding ability, we calculate the coverage degree (n-gram recall) between the generations and the input questions, as illustrated in the right part of Figure 4. The results indicate that Re2 increases the n-gram (n=1,2,3,4) recall in the output explanations, underscoring how our method enhances the model’s focus on the question during the reasoning process.
Prompt Vanilla CoT
P0
Q: {question}
#Answer format instruction#
A: Let’s think step by step.
77.79 78.77
P1
Q: {question}
Read the question again: {question}
#Answer format instruction#
A: Let’s think step by step.
79.45 80.59
P2
Q: {question}
Q: {question}
#Answer format instruction#
A: Let’s think step by step.
78.09 79.38
P3
Q: {question}
A: Let’s think step by step.
Read the question again: {question}
#Answer format instruction#
A: Let’s think step by step.
79.08 80.36
P4
Q: {question}
A: Let’s think step by step.
Q: {question}
#Answer format instruction#
A: Let’s think step by step.
78.09 79.38
Table 7: Results of different re-reading instructions.
The impact of different re-reading instructions
We further conduct experiments to examine the influence of Re2 within the context of CoT prompting. Specifically, we design various instructions for question re-reading using ChatGPT on GSM8K dataset. As depicted in P1 and P2 in Table 7, instruction P1, which includes the phrase ‘‘Read the question again:", exhibits superior performance compared to directly repeating the question twice. These results suggest that providing more detailed re-reading instructions to the LLMs is advantageous. Subsequently, we explore the possibility of introducing re-reading for CoT instructions (i.e. repeating ‘‘Let’s think step by step"), as exemplified in P3 and P4. However, we observe that repeating the thinking process two times does not yield any discernible benefits. It’s noteworthy that, in general, question re-reading consistently improves reasoning performance compared to the standard CoT prompting without question re-reading (P0).
4 Related Work
Reasoning with Large Language Models.
LLMs represent a significant milestone in the journey towards artificial general intelligence (AGI) OpenAI (2023); Touvron et al. (2023b) . Reasoning ability is particularly crucial on the way towards AGI, where artificial intelligence needs to act or think like human beings Qiao et al. (2023); Huang and Chang (2023). In the literature on LLMs, performing reasoning tasks via interaction in natural language plays a significant role in evaluating an LLM, into which academia and industry have been dedicating many endeavors Wei et al. (2022a); Suzgun et al. (2023b); Turpin et al. (2023). In principle, most works for reasoning with large language models could fall into the paradigm of “Chain-of-Though” Wei et al. (2022b); Kojima et al. (2022), which assists LLMs in fulfilling complex reasoning tasks by generating intermediate steps explicitly. Therefore, most of the endeavors are dedicated to improving the basic principle by the following aspects: i) the structure of “chain”, e.g., tree Yao et al. (2023a), graph Yao et al. (2023b); ii) the modality of the chain, e.g., program Gao et al. (2023); iii) the reliability of the chain, e.g., self-consistency Wang et al. (2023c), faithful Lyu et al. (2023), retrieval-based verifying He et al. (2023); and iv) decomposition of the chain, e.g., least-to-most Zhou et al. (2023), decomposed Radhakrishnan et al. (2023), plan-to-solve Wang et al. (2023a). In contrast, our simple re-reading strategy for LLMs is orthogonal to these improvements via a trade-off between the intermediate steps and the query itself. Besides, our re-reading strategy is complementary to many previous works by preventing the answer from being derived overwhelmingly from the CoT but overlooking the original query.
Re-reading Strategy in Text Understanding.
In deep learning, the success of performing text-understanding tasks Song et al. (2018); Luo et al. (2019a); Yang et al. (2019); Lei et al. (2019) depends on the heuristics of human reading strategy, e.g., pre-reading, re-reading and post-reading Saricoban (2002); Toprak and ALMACIOĞLU (2009); Pressley and Afflerbach (2012); Ozek and Civelek (2006); Dowhower (1989). Specifically, many effective algorithms have been crafted around the idea of re-reading. Although deep architectures, from multi-layer Bi-LSTM Huang et al. (2015) to Transformer-encoder Vaswani et al. (2017), have their mechanisms that provide a form of “re-reading”, the notion that simply processing an input once might not be sufficient for understanding or generating a complex output has been long-standing. Initially, Sha et al. (2016) and Sha et al. (2017) found that repeated reading mechanisms do improve performance on some tasks, e.g., sentiment analysis, semantic relation classification, and event extraction. Then, Liu and Li (2016) propose to mimic the repeated reading strategy and present neural networks with multi-level attention, which is proven effective in recognizing implicit discourse relations. Sequentially, Zhu et al. (2018) propose a multi-glance mechanism, modeling the habit of reading behavior, which can benefit a wide range of tasks. More recently, Luo et al. (2019b) adopt a network to encode the gist of paragraphs for rough reading and a decision-making policy for careful reading, which can improve extractive summarization. Therefore, it is natural to introduce a re-reading strategy to LLMs since the Transformer-decoder architecture of LLMs, with unidirectional attention mechanisms, hinders the implicit re-reading capability.
Knowledge Recall.
From the perspective of information seeking, prompting LLMs can be seen as a sort of “knowledge recall” via a parametric fashion, where the prompt can be seen as a retrieval query. In contrast to conventional non-parametric retrieval – vector database Karpukhin et al. (2020); Izacard et al. (2022) for example, the LLM as a neural knowledge model Bosselut et al. (2019); AlKhamissi et al. (2022) can easily generalize for huge knowledge coverage, contributing to its efficacy in broad applications. In the context of CoT reasoning, Chen et al. (2023) conjuncture that LLM can be exposed to certain CoTs during training and easily complete reasoning by knowledge recall. As such, it is natural to adapt the basic but prevalent query augmentation technique in the term-based retrieval domain Dai and Callan (2019), which repeats the original query multiple times over the augmented part Wang et al. (2023b); Shen et al. (2023), into prompting LLMs.
5 Conclusion and Future Works
This paper introduces Re2, a simple and effective prompting method for LLM reasoning that improves performance by "re-reading" the question. By shifting focus to the input phase, Re2 operates independently from other thought-eliciting promptings. Moreover, it shows promise in fostering bidirectional comprehension of questions in decoder-only LLMs. Our comprehensive experiments cover a wide range of reasoning benchmarks, diverse LLM types, various task settings, and compatibility assessments with other prompting methods, validating the efficacy and versatility of Re2. Our findings encourage the research community to prioritize a deeper understanding of input questions, thereby complementing existing thought-eliciting prompting strategies. Future endeavors will aim to explore its versatility in additional contexts beyond reasoning, including multi-turn dialogue and multi-modal reasoning applications.
6 Limitations
In this paper, we introduce a simple yet effective prompting method for enhancing reasoning in LLMs and conduct extensive experiments to validate its effectiveness. Despite our best efforts, there may be still some limitations that remain in our study. Our investigation primarily revolves around empirical studies with extensive experiments to validate Re2, similar to most works in prompting research Zheng et al. (2023); Yin et al. (2023); Gao et al. (2023). Future efforts will include more theoretical analyses to provide a solid foundation. Additionally, Re2 marginally increases the input length, potentially leading to a slight reduction in efficiency for longer questions during inference. Furthermore, our paper concentrates exclusively on the impact of Re2 within the reasoning domain. Future work will extend the exploration to additional contexts, such as multi-turn dialogue and multi-modal reasoning.
7 Ethics
We conducted experiments on seven mathematical reasoning benchmarks, comprising GSM8K Cobbe et al. (2021), SVAMP Patel et al. (2021), ASDiv Miao et al. (2020), AQuA Ling et al. (2017), AddSub Hosseini et al. (2014), MultiArith Roy and Roth (2015), SingelEQ Roy et al. (2015), three commonsense reasoning benchmarks (CSQA Talmor et al. (2019), StrategyQA Geva et al. (2021), and ARC Clark et al. (2018)), and two symbolic benchmarks (Date Understanding Suzgun et al. (2023a) and Coinflip Wei et al. (2022b)). Among these, GSM8K and SVAMP datasets utilize code under the MIT License, while AQuA, StrategyQA, Date Understanding, Coinflip utilize code under the Apache-2.0 license, and ARC utilizes code under CC-BY-SA-4.0. The licenses for the remaining datasets are unspecified.
The proposed prompts do not involve the collection or utilization of personal information pertaining to other individuals. Details regarding the prompts used in our experiments are provided in Appendix §B. Furthermore, it is ensured that the prompts utilized in this research do not pose any threat to the safety or well-being of others.
Appendix A Datasets
Table 8 presents statistics and examples for the reasoning benchmarks we considered.
Appendix B Specific Prompting Methods
Detailed information regarding various promptings is shown in Table 9 and Table 10. The instructions of answer-format can be found in Table 11.
Appendix C Attention Analysis
Figure 5: Attention visualization with and without Re2. (a) CoT prompting: there is only one pass for the question. (b) CoT+Re2 re-reads the question, including first pass and second pass. The row of matrix represents the query tokens and the column represents the key tokens.
To gain deeper insights into how Re2 reshapes attention during inference, we visualize the attention distribution by computing the average attention weights across all heads and layers in Llama-2. The results are presented in Figure 5, revealing two key findings: (1) In the block of “Second pass” attending to the “First pass” as shown in (b) for CoT+Re2, we observe explicit attentions in the upper triangle. This observation indicates that tokens in the second question can focus on the tokens behind the corresponding positions in the first question. In this way, Re2 enables a “bidirectional” understanding of the question. Notably, with the inclusion of Re2, the generation process maintains a higher attention weight on the question tokens. By calculating the proportion of attention weights assigned to the question tokens during generation, we observe an increase from 0.32 to 0.40 with the utilization of Re2. This finding suggests that the re-reading mechanism enhances the model’s focus on the question during the reasoning process.
Appendix D Case Study
We also conduct a case study to show the effectiveness of our proposed re-reading prompting over the chain-of-thought. We choose two examples from GSM, and the results generated by ChatGPT are listed in Table  12-13. It is evident that our method can better align the evidence in the question with the corresponding explanation hints. We can observe that CoT+Re2 tends to highlight the important evidences in the question before generating the explanation, for example, “In the morning, she gives 15 cups of feed, and in the afternoon, she gives another 25. So …" in Table 12 and “The bonus is worth half a month’s salary, which is …" in Table 13. This observation is also consistent with the right figure in Figure 4.
Appendix E More Cases
Tables 14-18 provide more examples generated by ChatGPT with CoT and CoT+Re2. We also provide several examples generated by davinci-003 and ChatGPT in the Vanilla prompting (e.g. no instruction) in Tables 19-22. They show that ChatGPT with Vanilla directly generates answer in Coin Filp and Date Understanding dataset (Tables 19-20), but still generates CoT output in some other datasets (Tables 21-22).
Methods Prompt Content
Vanilla
Q: {question}
#Answer format instruction#
A:
Vanilla+Re2
Q: {question}
Read the question again: {question}
#Answer format instruction#
A:
CoT
Q: {question}
#Answer format instruction#
A: Let’s think step by step.
CoT+Re2
Q: {question}
Read the question again: {question}
#Answer format instruction#
A: Let’s think step by step.
PS
Q: {question}
#Answer format instruction#
A: Let’s first understand the problem and devise a plan to solve the problem.
Then, let’s carry out the plan, solve the problem step by step,
and give the ultimate answer. Please explicitly generate the mentioned process:
[Problem Understanding], [Plan], [Solving/Calculations], [Answer]. in your response.
PS+Re2
Q: {question}
Read the question again: {question}
#Answer format instruction#
A: Let’s first understand the problem and devise a plan to solve the problem.
Then, let’s carry out the plan, solve the problem step by step,
and give the ultimate answer. Please explicitly generate the mentioned process:
[Problem Understanding], [Plan], [Solving/Calculations], [Answer]. in your response.
Table 9: Specific prompts of Vanilla, Vanilla+Re2, CoT, CoT+Re2, PS, and PS+Re2.
Methods Prompt Content
PAL
#!/bin/python3
import math
import numpy as np
import statistics
import sympy as sp
########### Task Instruction ##############
# You will write python program to solve math problems.
# You will only write code blocks.
# Please generate your code block in ‘def solution()‘ function, and thus it can be
executed by python interpreter. You don’t need to call ‘solution()‘ function because
it will be called by the system.
# The concrete format of ‘solution()‘ is as follows:
# def solution():
# """<question>"""
# <your code>
# result = <your result>
# return result
# Q: {question}
# Your defined "solution()" function with comments here.
PAL+Re2
#!/bin/python3
import math
import numpy as np
import statistics
import sympy as sp
########### Task Instruction ##############
# You will write python program to solve math problems.
# You will only write code blocks.
# Please generate your code block in ‘def solution()‘ function, and thus it can be
executed by python interpreter. You don’t need to call ‘solution()‘ function because
it will be called by the system.
# The concrete format of ‘solution()‘ is as follows:
# def solution():
# """<question>"""
# <your code>
# result = <your result>
# return result
# Q: {question}
# Read the question again: {question}
# Your defined "solution()" function with comments here.
Table 10: Specific prompts of PAL and PAL+Re2
Tasks Answer-format Instructions
GSM
Your final answer should be a single numerical number, in the form $𝑎𝑛𝑠𝑤𝑒𝑟$,
at the end of your response.
SVAMP
Your final answer should be a single numerical number, in the form $𝑎𝑛𝑠𝑤𝑒𝑟$,
at the end of your response.
ASDIV
Your final answer should be a single numerical number, in the form $𝑎𝑛𝑠𝑤𝑒𝑟$,
at the end of your response.
AQUA
[…]
MultiArith
Your final answer should be a single numerical number, in the form $𝑎𝑛𝑠𝑤𝑒𝑟$,
at the end of your response.
SingleEQ
Your final answer should be a single numerical number, in the form $𝑎𝑛𝑠𝑤𝑒𝑟$,
at the end of your response.
AddSub
Your final answer should be a single numerical number, in the form $𝑎𝑛𝑠𝑤𝑒𝑟$,
at the end of your response.
CommonsenseQA
[…]
StrategyQA
Your answer should be either $𝑦𝑒𝑠$ or $𝑛𝑜$, in the form $𝑎𝑛𝑠𝑤𝑒𝑟$.
ARC
[…]
Date Understanding
Your answer should be a date, in the format of $𝑀𝑀𝐷𝐷𝑌𝑌𝑌𝑌$,
e.g. 05/01/20220501202205 / 01 / 2022.
Coin Flip
Your answer should be either $𝑦𝑒𝑠$ or $𝑛𝑜$, in the form $𝑎𝑛𝑠𝑤𝑒𝑟$.
Table 11: Specific answer-format instruction in each task.
Q: Every day, Wendi feeds each of her chickens three cups of mixed chicken feed, containing seeds, mealworms and vegetables to help keep them healthy. She gives the chickens their feed in three separate meals. In the morning, she gives her flock of chickens 15 cups of feed. In the afternoon, she gives her chickens another 25 cups of feed. How many cups of feed does she need to give her chickens in the final meal of the day if the size of Wendi’s flock is 20 chickens? (Answer: 20)
CoT
Input:
Q: Every day, Wendi feeds each of her chickens three cups of mixed chicken feed, containing seeds, mealworms and vegetables to help keep them healthy. She gives the chickens their feed in three separate meals. In the morning, she gives her flock of chickens 15 cups of feed. In the afternoon, she gives her chickens another 25 cups of feed. How many cups of feed does she need to give her chickens in the final meal of the day if the size of Wendi’s flock is 20 chickens?
Your final answer should be a single numerical number, in the form $𝑎𝑛𝑠𝑤𝑒𝑟$, at the end of your response.
A: Let’s think step by step.
Output:
First, let’s calculate how much feed Wendi gives her chickens in the morning. She has 20 chickens, and she gives each chicken 3 cups of feed, so in total, she gives them 20 x 3 = 60 cups of feed in the morning. Next, let’s calculate how much feed Wendi gives her chickens in the afternoon. Again, she has 20 chickens, and she gives each chicken 3 cups of feed, so in total, she gives them 20 x 3 = 60 cups of feed in the afternoon. To find out how much feed she needs to give her chickens in the final meal of the day, we need to subtract the total feed given in the morning and afternoon from the total feed needed for the day. The total feed needed for the day for 20 chickens is 20 x 3 x 3 = 180 cups of feed. So, the feed needed in the final meal of the day is 180 - 60 - 60 = 60 cups of feed. Therefore, Wendi needs to give her chickens 60 cups of feed in the final meal of the day. Answer: 606060
CoT+Re2
Input:
Q: Every day, Wendi feeds each of her chickens three cups of mixed chicken feed, containing seeds, mealworms and vegetables to help keep them healthy. She gives the chickens their feed in three separate meals. In the morning, she gives her flock of chickens 15 cups of feed. In the afternoon, she gives her chickens another 25 cups of feed. How many cups of feed does she need to give her chickens in the final meal of the day if the size of Wendi’s flock is 20 chickens?
Read the question again: Every day, Wendi feeds each of her chickens three cups of mixed chicken feed, containing seeds, mealworms and vegetables to help keep them healthy. She gives the chickens their feed in three separate meals. In the morning, she gives her flock of chickens 15 cups of feed. In the afternoon, she gives her chickens another 25 cups of feed. How many cups of feed does she need to give her chickens in the final meal of the day if the size of Wendi’s flock is 20 chickens?
Your final answer should be a single numerical number, in the form $𝑎𝑛𝑠𝑤𝑒𝑟$, at the end of your response.
A: Let’s think step by step.
Output:
First, we need to determine the total amount of feed Wendi gives her chickens in the morning and afternoon. In the morning, she gives 15 cups of feed, and in the afternoon, she gives another 25 cups. So, the total amount of feed given in the morning and afternoon is 15 + 25 = 40 cups. Next, we need to determine the total amount of feed needed for Wendi’s flock of 20 chickens. Since each chicken requires 3 cups of feed, we multiply 3 by 20 to get 60 cups. Finally, we subtract the total amount of feed given in the morning and afternoon from the total amount needed. 60 - 40 = 20 Therefore, Wendi needs to give her chickens 20 cups of feed in the final meal of the day. Answer: 202020
Table 12: GSM Example 4 (Complexity=3)
Q: Tim gets a promotion that offers him a 5% raise on his \$20000 a month salary. It also gives him a bonus worth half a month’s salary. How much money will he make in a year? (Answer: 262500)
CoT
Input:
Q: Tim gets a promotion that offers him a 5% raise on his \$20000 a month salary. It also gives him a bonus worth half a month’s salary. How much money will he make in a year?
Your final answer should be a single numerical number, in the form $𝑎𝑛𝑠𝑤𝑒𝑟$, at the end of your response.
A: Let’s think step by step.
Output:
First, let’s calculate the 5% raise on Tim’s \$20000 monthly salary. To do this, we multiply \$20000 by 5% (or 0.05):
\$20000 * 0.05 = \$1000
So, Tim’s raise is \$1000 per month.
Next, let’s calculate the bonus worth half a month’s salary. Half of \$20000 is \$10000, so Tim’s bonus is \$10000.
Now, let’s calculate Tim’s total monthly salary after the raise and bonus:
\$20000 (original salary) + \$1000 (raise) + \$10000 (bonus) = \$31000
Finally, let’s calculate Tim’s annual salary by multiplying his monthly salary by 12:
\$31000 * 12 = \$372000
Therefore, Tim will make \$372000 in a year. So, the answer is 372000372000372000.
CoT+Re2
Input:
Q: Tim gets a promotion that offers him a 5% raise on his \$20000 a month salary. It also gives him a bonus worth half a month’s salary. How much money will he make in a year?
Read the question again: Tim gets a promotion that offers him a 5% raise on his \$20000 a month salary. It also gives him a bonus worth half a month’s salary. How much money will he make in a year?
Your final answer should be a single numerical number, in the form $𝑎𝑛𝑠𝑤𝑒𝑟$, at the end of your response.
A: Let’s think step by step.
Output:
First, let’s calculate the 5% raise on Tim’s \$20000 monthly salary.
The raise amount would be 5% of \$20000, which is equal to 0.05 * \$20000 = \$1000.
So, Tim’s new monthly salary would be \$20000 + \$1000 = \$21000.
Next, let’s calculate the bonus amount. The bonus is worth half a month’s salary, which is equal to 0.5 * \$21000 = \$10500.
Now, to find out how much money Tim will make in a year, we need to calculate the total salary for 12 months and add the bonus.
Total salary for 12 months = \$21000 * 12 = \$252000.
So, Tim will make a total of \$252000 + \$10500 = \$262500 in a year.
Therefore, Tim will make 262500262500262500 in a year.
Table 13: GSM Example 641 (Complexity=6)
Q: Colby loves going to the movies and every month his parents give him \$150 to spend at the movies. Tickets for Fridays and Saturdays cost \$10. Tickets for any other day cost \$7. Popcorn costs \$8 and boxes of candy cost \$2. It is the last day of the month and it’s a Friday. He wants to make sure he gets a popcorn and box of candy that night. How many movies can he see if he already saw 5 movies on a Friday or Saturday, 8 movies on other days, had 2 tubs of popcorn, and four boxes of candy that month? (Answer: 1)
CoT
Input:
Q: Colby loves going to the movies and every month his parents give him \$150 to spend at the movies. Tickets for Fridays and Saturdays cost \$10. Tickets for any other day cost \$7. Popcorn costs \$8 and boxes of candy cost \$2. It is the last day of the month and it’s a Friday. He wants to make sure he gets a popcorn and box of candy that night. How many movies can he see if he already saw 5 movies on a Friday or Saturday, 8 movies on other days, had 2 tubs of popcorn, and four boxes of candy that month?
Your final answer should be a single numerical number, in the form $𝑎𝑛𝑠𝑤𝑒𝑟$, at the end of your response.
A: Let’s think step by step.
Output:
First, let’s calculate the cost of the movies Colby has already seen. He has seen 5 movies on a Friday or Saturday, which cost \$10 each, so that is a total of 5 * \$10 = \$50. He has also seen 8 movies on other days, which cost \$7 each, so that is a total of 8 * \$7 = \$56. Therefore, the total cost of the movies he has already seen is \$50 + \$56 = \$106. Next, let’s calculate the cost of the popcorn and candy he wants to buy on the last day of the month. The popcorn costs \$8 and the candy costs \$2, so the total cost of the snacks is \$8 + \$2 = \$10. Now, let’s subtract the cost of the movies Colby has already seen and the cost of the snacks from the \$150 his parents gave him. \$150 - \$106 - \$10 = \$34. Since each movie on a Friday or Saturday costs \$10 and Colby has \$34 left, he can see a maximum of \$34 / \$10 = 3 movies on a Friday or Saturday. In conclusion, Colby can see a total of 3 + 5 = 8 movies in the month. 888
CoT+Re2
Input:
Q: Colby loves going to the movies and every month his parents give him \$150 to spend at the movies. Tickets for Fridays and Saturdays cost \$10. Tickets for any other day cost \$7. Popcorn costs \$8 and boxes of candy cost \$2. It is the last day of the month and it’s a Friday. He wants to make sure he gets a popcorn and box of candy that night. How many movies can he see if he already saw 5 movies on a Friday or Saturday, 8 movies on other days, had 2 tubs of popcorn, and four boxes of candy that month?
Read the question again: Colby loves going to the movies and every month his parents give him \$150 to spend at the movies. Tickets for Fridays and Saturdays cost \$10. Tickets for any other day cost \$7. Popcorn costs \$8 and boxes of candy cost \$2. It is the last day of the month and it’s a Friday. He wants to make sure he gets a popcorn and box of candy that night. How many movies can he see if he already saw 5 movies on a Friday or Saturday, 8 movies on other days, had 2 tubs of popcorn, and four boxes of candy that month?
Your final answer should be a single numerical number, in the form $𝑎𝑛𝑠𝑤𝑒𝑟$, at the end of your response.
A: Let’s think step by step.
Output:
First, let’s calculate the total amount Colby has spent so far: For the 5 movies on Fridays and Saturdays, he spent 5 x \$10 = \$50. For the 8 movies on other days, he spent 8 x \$7 = \$56. For the 2 tubs of popcorn, he spent 2 x \$8 = \$16. For the 4 boxes of candy, he spent 4 x \$2 = \$8. Adding all these expenses together, Colby has spent \$50 + \$56 + \$16 + \$8 = \$130. Now, let’s calculate how much money he has left from the \$150 his parents gave him: Colby started with \$150 - \$130 = \$20. Since it’s the last day of the month and it’s a Friday, he wants to buy a popcorn (\$8) and a box of candy (\$2). This means he will spend \$8 + \$2 = \$10. Therefore, Colby will have \$20 - \$10 = \$10 left. Now, let’s calculate how many movies he can see with the remaining \$10: Since tickets on Fridays and Saturdays cost \$10, and he has \$10 left, he can see 1 movie on a Friday or Saturday. Therefore, Colby can see 1 more movie on the last day of the month. So, the final answer is 111.
Table 14: GSM Example 926 (Complexity>=9)
Q: Mike was a pen pal with 5 people. He stopped being penpals with 2 of them. They each send 2 letters a week that are 5 pages long. He responds in kind. He can write a page every 6 minutes. How many hours does he spend writing a week? (Answer: 3)
CoT
Input:
Q: Mike was a pen pal with 5 people. He stopped being penpals with 2 of them. They each send 2 letters a week that are 5 pages long. He responds in kind. He can write a page every 6 minutes. How many hours does he spend writing a week?
Your final answer should be a single numerical number, in the form $𝑎𝑛𝑠𝑤𝑒𝑟$, at the end of your response.
A: Let’s think step by step.
Output:
First, let’s calculate how many letters Mike receives in a week from the 5 pen pals. Each pen pal sends 2 letters per week, so 5 pen pals would send him 5 * 2 = 10 letters in a week. Next, let’s calculate how many pages Mike needs to write in response to each letter. Each letter is 5 pages long, and Mike responds in kind, so he would write 5 pages for each letter he receives. Therefore, Mike needs to write a total of 10 * 5 = 50 pages in a week. Now, let’s calculate how many minutes it takes for Mike to write one page. We are given that he can write a page in 6 minutes. Since Mike needs to write 50 pages, it will take him 50 * 6 = 300 minutes to complete all the writing. Lastly, let’s convert the minutes to hours. There are 60 minutes in an hour, so 300 minutes is equal to 300 / 60 = 5 hours. Therefore, Mike spends 5 hours writing in a week. Answer: 555.
CoT+Re2
Input:
Q: Mike was a pen pal with 5 people. He stopped being penpals with 2 of them. They each send 2 letters a week that are 5 pages long. He responds in kind. He can write a page every 6 minutes. How many hours does he spend writing a week?
Read the question again: Mike was a pen pal with 5 people. He stopped being penpals with 2 of them. They each send 2 letters a week that are 5 pages long. He responds in kind. He can write a page every 6 minutes. How many hours does he spend writing a week?
Your final answer should be a single numerical number, in the form $𝑎𝑛𝑠𝑤𝑒𝑟$, at the end of your response.
A: Let’s think step by step.
Output:
First, let’s calculate the total number of letters that Mike receives in a week. Since he was pen pals with 5 people, and he stopped being pen pals with 2 of them, he is still receiving letters from 5 - 2 = 3 people. Each of these 3 people sends 2 letters a week, so the total number of letters that Mike receives in a week is 3 * 2 = 6 letters. Next, let’s calculate the total number of pages that Mike needs to write in response. Each letter is 5 pages long, so the total number of pages that Mike needs to write is 6 * 5 = 30 pages. Since Mike can write a page every 6 minutes, the total time he spends writing in a week is 30 * 6 = 180 minutes. To convert this to hours, we divide by 60 since there are 60 minutes in an hour. So, Mike spends 180 / 60 = 3 hours writing in a week. Therefore, Mike spends 333 hours writing a week.
Table 15: GSM Example 242 (Complexity=6)
Q: Terri is knitting a sweater with two sleeves, a collar, and a decorative rosette. The body of the sweater takes 900 stitches to complete, the collar takes a tenth of that number of stitches, and the rosette takes twice as many as the collar. The whole sweater is an 1800-stitch project. How many stitches does each sleeve take? (Answer: 315)
CoT
Input:
Q: Terri is knitting a sweater with two sleeves, a collar, and a decorative rosette. The body of the sweater takes 900 stitches to complete, the collar takes a tenth of that number of stitches, and the rosette takes twice as many as the collar. The whole sweater is an 1800-stitch project. How many stitches does each sleeve take?
Your final answer should be a single numerical number, in the form $𝑎𝑛𝑠𝑤𝑒𝑟$, at the end of your response.
A: Let’s think step by step.
Output:
First, we know that the body of the sweater takes 900 stitches to complete. Next, we know that the collar takes a tenth of the number of stitches as the body. So, the collar takes 900 stitches / 10 = 90 stitches. Finally, we know that the rosette takes twice as many stitches as the collar. So, the rosette takes 90 stitches * 2 = 180 stitches. To find out how many stitches each sleeve takes, we need to subtract the stitches of the body, collar, and rosette from the total number of stitches in the sweater. So, each sleeve takes 1800 stitches - 900 stitches (body) - 90 stitches (collar) - 180 stitches (rosette) = 630 stitches. Therefore, each sleeve takes 630630630 stitches.
CoT+Re2
Input:
Q: Terri is knitting a sweater with two sleeves, a collar, and a decorative rosette. The body of the sweater takes 900 stitches to complete, the collar takes a tenth of that number of stitches, and the rosette takes twice as many as the collar. The whole sweater is an 1800-stitch project. How many stitches does each sleeve take?
Read the question again: Terri is knitting a sweater with two sleeves, a collar, and a decorative rosette. The body of the sweater takes 900 stitches to complete, the collar takes a tenth of that number of stitches, and the rosette takes twice as many as the collar. The whole sweater is an 1800-stitch project. How many stitches does each sleeve take?
Your final answer should be a single numerical number, in the form $𝑎𝑛𝑠𝑤𝑒𝑟$, at the end of your response.
A: Let’s think step by step.
Output:
First, we know that the body of the sweater takes 900 stitches to complete. Next, we are told that the collar takes a tenth of the number of stitches in the body of the sweater. Since the body has 900 stitches, the collar will have 900 / 10 = 90 stitches. Finally, we are told that the rosette takes twice as many stitches as the collar. Since the collar has 90 stitches, the rosette will have 90 * 2 = 180 stitches. Now, let’s add up all the stitches to find the total number of stitches in the sweater: 900 (body) + 90 (collar) + 180 (rosette) = 1170 stitches. We are given that the whole sweater is an 1800-stitch project. So, the remaining stitches must be for the sleeves. 1800 (total stitches) - 1170 (stitches accounted for) = 630 stitches for the sleeves. Since there are two sleeves, each sleeve will have 630 / 2 = 315 stitches. Therefore, each sleeve takes 315 stitches.
Table 16: GSM Example 273 (Complexity=6)
Q: Solo has to read 4 pages from his Science textbook, 20 pages from his Social Studies textbook, 7 pages from his History textbook and 8 pages from his Geography textbook. Solo read 15 pages on Monday. If he has 4 more days to complete all his reading, how many pages does he need to read, on average, in one day? (Answer: 6)
CoT
Input:
Q: Solo has to read 4 pages from his Science textbook, 20 pages from his Social Studies textbook, 7 pages from his History textbook and 8 pages from his Geography textbook. Solo read 15 pages on Monday. If he has 4 more days to complete all his reading, how many pages does he need to read, on average, in one day?
Your final answer should be a single numerical number, in the form $𝑎𝑛𝑠𝑤𝑒𝑟$, at the end of your response.
A: Let’s think step by step.
Output:
First, let’s calculate how many pages Solo has left to read. He needs to read 4 pages from his Science textbook, 20 pages from his Social Studies textbook, 7 pages from his History textbook, and 8 pages from his Geography textbook. So, the total number of pages he needs to read is 4 + 20 + 7 + 8 = 39 pages. Since Solo has 4 more days to complete all his reading, we can divide the total number of pages by the number of days. Therefore, Solo needs to read an average of 39 pages / 4 days = 9.75 pages per day. Therefore, Solo needs to read an average of 9.759.759.75 pages per day.
CoT+Re2
Input:
Q: Solo has to read 4 pages from his Science textbook, 20 pages from his Social Studies textbook, 7 pages from his History textbook and 8 pages from his Geography textbook. Solo read 15 pages on Monday. If he has 4 more days to complete all his reading, how many pages does he need to read, on average, in one day?
Read the question again: Solo has to read 4 pages from his Science textbook, 20 pages from his Social Studies textbook, 7 pages from his History textbook and 8 pages from his Geography textbook. Solo read 15 pages on Monday. If he has 4 more days to complete all his reading, how many pages does he need to read, on average, in one day?
Your final answer should be a single numerical number, in the form $𝑎𝑛𝑠𝑤𝑒𝑟$, at the end of your response.
A: Let’s think step by step.
Output:
First, let’s calculate the total number of pages Solo needs to read. He needs to read 4 pages from his Science textbook, 20 pages from his Social Studies textbook, 7 pages from his History textbook, and 8 pages from his Geography textbook. So, the total number of pages he needs to read is 4 + 20 + 7 + 8 = 39 pages. Next, let’s calculate the number of pages he has left to read after Monday. He read 15 pages on Monday, so he has 39 - 15 = 24 pages left to read. Finally, let’s calculate the average number of pages he needs to read per day. He has 4 more days to complete his reading, so he needs to read 24 / 4 = 6 pages per day. Therefore, Solo needs to read an average of 666 pages per day.
Table 17: GSM Example 101 (Complexity=6)
Q: Would a nickel fit inside a koala pouch? (Answer: yes)
CoT
Input:
Q: Would a nickel fit inside a koala pouch?
Your answer should be either $𝑦𝑒𝑠$ or $𝑛𝑜$, in the form $𝑎𝑛𝑠𝑤𝑒𝑟$.
A: Let’s think step by step.
Output:
First, let’s consider the size of a nickel. A nickel has a diameter of approximately 21.21 mm and a thickness of approximately 1.95 mm. Next, let’s consider the size of a koala pouch. A koala pouch is designed to hold a baby koala, also known as a joey. The size of a koala pouch can vary, but it is generally not large enough to fit objects the size of a nickel. Therefore, based on the size of a nickel and the size of a koala pouch, it is unlikely that a nickel would fit inside a koala pouch. $𝑁𝑜$
CoT+Re2
Input:
Q: Would a nickel fit inside a koala pouch?
Read the question again: Would a nickel fit inside a koala pouch?
Your answer should be either $𝑦𝑒𝑠$ or $𝑛𝑜$, in the form $𝑎𝑛𝑠𝑤𝑒𝑟$.
A: Let’s think step by step.
Output:
First, let’s consider the size of a nickel. A nickel has a diameter of about 0.835 inches (21.21 mm) and a thickness of about 0.077 inches (1.95 mm). Next, let’s consider the size of a koala pouch. A koala pouch is designed to hold a baby koala, also known as a joey. The pouch is stretchy and expands as the joey grows. However, the size of the pouch may vary depending on the age and size of the joey. Based on the size of a nickel and the stretchiness of a koala pouch, it is possible that a nickel could fit inside a koala pouch. However, it would depend on the size of the pouch and the position of the joey inside. So, the answer is $𝑦𝑒𝑠$, a nickel could potentially fit inside a koala pouch.
Table 18: StrategyQA Example 262
Input:
A coin is heads up. Hank does not flip the coin. Janine does not flip the coin. Frankie flips the coin. Isa flips the coin. Is the coin still heads up? Note that "flip" here means "reverse".
Your answer should be either $𝑦𝑒𝑠$ or $𝑛𝑜$, in the form $𝑎𝑛𝑠𝑤𝑒𝑟$.
A:
davinci-003 with Vanilla
Output:
No.
ChatGPT with Vanilla
Output:
No.
Table 19: Coin Flip Example 16. Without explicit instruction, ChatGPT and davinci-003 directly generate the final answer.
Input:
Yesterday was April 30, 2021. What is the date tomorrow in MM/DD/YYYY?
Your answer should be a date, in the format of $𝑀𝑀𝐷𝐷𝑌𝑌𝑌𝑌$, e.g. 05/01/20220501202205 / 01 / 2022.
A:
davinci-003 with Vanilla
Output:
05/01/2021
ChatGPT with Vanilla
Output:
11/13/2019.
Table 20: Date Understanding Example 1. Without explicit instruction, ChatGPT and davinci-003 directly generate the final answer.
Input:
Ned had to wash 9 short sleeve shirts and 21 long sleeve shirts before school. If he had only washed 29 of them by the time school started, how many did he not wash?
Your final answer should be a single numerical number, in the form $𝑎𝑛𝑠𝑤𝑒𝑟$, at the end of your response.
A:
davinci-003 with Vanilla
Output:
He did not wash 20 shirts. 202020
ChatGPT with Vanilla
Output:
Ned had to wash a total of 9 + 21 = 30 shirts.He washed 29 shirts, so he did not wash 30 - 29 = <<30-29=1>>1 shirt. Therefore, he did not wash 111 shirt.
Table 21: MultiArith Example 9. Without explicit instruction, ChatGPT still can generate the CoT output, but davinci-003 does not.
Input:
A trader sold an article at a profit of 20% for Rs.360. What is the cost price of the article? Answer Choices: A) 270, B) 300, C) 280, D) 320, E) 315
[…]
A:
davinci-003 with Vanilla
Output:
$𝐵$
ChatGPT with Vanilla
Output:
Let’s assume the cost price of the article is x. The selling price of the article is given as Rs. 360, which is 120% of the cost price (100% + 20% profit). So, we can write the equation as: x + 20% of x = 360 Simplifying the equation, we get: 1.2x = 360 Dividing both sides by 1.2, we find: x = 300 Therefore, the cost price of the article is Rs. 300. The correct answer is ($𝐵300$.
Table 22: AQUA Example 4. Without explicit instruction, ChatGPT still can generate the CoT output, but davinci- 003 does not.